OpenStudy (anonymous):

Which of the following probabilities is the greatest for a standard normal distribution? A. P(-1.5≤z≤-0.5) B. P(-0.5≤z≤0.5) C. P(0.5≤z≤1.5) D. P(1.5≤z≤2.5)

OpenStudy (kropot72):

This can be found by using a standard normal distribution table, such as the table here: http://www.math.bgu.ac.il/~ngur/Teaching/probability/normal.pdf

OpenStudy (kropot72):

If we take option A, from the table we find: cumulative probability for z = -1.5 is 0.0668 cumulative probability for z = -0.5 is 0.3085 Therefore P(-1.5≤z≤-0.5) = 0.3085 - 0.0668 = 0.2417 Now you need to repeat this process for the other answer options to find the greatest probability.

OpenStudy (anonymous):

Ohhh, I see. How would I find it for 2.5?

OpenStudy (kropot72):

Just look it up in the table that I gave you a link to.

OpenStudy (anonymous):

Would it be B?

OpenStudy (kropot72):

You need to calculate the probability for each of the options.

OpenStudy (anonymous):

B came out to be 0.2909. I just want to know if I'm right?

OpenStudy (kropot72):

Your calculation for B is not correct. cumulative probability for z = 0.5 is 0.6915 cumulative probability for z = -0.5 is 0.3085 Therefore P(-0.5≤z≤0.5) = 0.6915 - 0.3085 = 0.383